# Physics-Informed LLM-Agent for Automated Modulation Design in Power Electronics Systems

### v1

@@ Abstract.
LLM-based autonomous agents have demonstrated outstanding performance in solving complex industrial tasks. However, in the pursuit of carbon neutrality and high-performance renewable energy systems, existing AI-assisted design automation faces significant limitations in explainability, scalability, and usability. To address these challenges, we propose LP-COMDA, an LLM-based, physics-informed autonomous agent that automates the modulation design of power converters in Power Electronics Systems with minimal human supervision. Unlike traditional AI-assisted approaches, LP-COMDA contains an LLM-based planner that gathers and validates design specifications through a user-friendly chat interface. The planner then coordinates with physics-informed design and optimization tools to iteratively generate and refine modulation designs autonomously. Through the chat interface, LP-COMDA provides an explainable design process, presenting explanations and charts. Experiments show that LP-COMDA outperforms all baseline methods, achieving a 63.2% reduction in error compared to the second-best benchmark method in terms of standard mean absolute error. Furthermore, empirical studies with 20 experts conclude that design time with LP-COMDA is over 33 times faster than conventional methods, showing its significant improvement on design efficiency over the current processes.
Agentic AI, PINN, LLM, Power Conversion
copyright: ccconference: Make sure to enter the correct conference title from your rights confirmation email; August 01–06, 2025; Toronto, Canadaccs: Information systems Language modelsccs: Computing methodologies Intelligent agentsccs: Hardware Power conversion
1. Introduction
@@
As temperature rises year on year due to the effect of Global Warming, there is an urgent need to reorient our economies towards carbon neutrality, especially in power generation. A significant penetration of renewable energy resources (RES) in power generation is crucial. Power Electronics Systems (PES) are indispensable for integrating RES into the contemporary power grid for three primary reasons (Lin, 2022). Firstly, power converters within PES transform the direct current (DC) output from solar panels or wind turbines to alternating current (AC), facilitating the seamless integration of RES into the AC grid. Secondly, PES manages power during peak and off-peak periods, guiding the charge and discharge processes of energy storage units. Lastly, it upholds grid stability through functions like voltage regulation and reactive power compensation. Achieving the desired outcomes necessitates meticulous modulation of power converters in PES, which involves controlling their switching to manage output voltage or current.
@@
The escalating adoption of RES and the burgeoning scale of multi-resource power systems complicate the modulation design for power converters within PES. The diversity of energy sources linked to the power grid intensifies system complexity. As more converter subsystems interlink, the modulation design problem’s dimensionality surges. Consequently, identifying an optimized design solution through manual computation becomes infeasible given this complexity. As cities worldwide embrace multi-resource power systems, modulation designs must be customized, catering to particular geographies and application objectives. Sometimes, these objectives can conflict, further complicating matters.
@@
The recent advancements in artificial intelligence (AI) and Large Language Models (LLMs) offer a potential resolution to these modulation design challenges through automation. Emerging research has introduced AI-centric modulation design techniques. For instance, recent works explore the use of an LLM with in-context learning for modulation design  (Lin et al., 2024). Extreme gradient boosting algorithm (XGBoost) was utilised to construct the surrogate model for the triple phase shift modulation strategy for the dual active bridge (DAB) converter using training data sourced from simulation tools or hardware prototypes (Lin et al., 2023a). Subsequently, a differential evolution algorithm collaborates iteratively with XGBoost until optimal modulation parameters are identified. Likewise, the Q-learning algorithm, a conventional reinforcement learning technique, can be trained offline to derive optimized modulation parameters (Tang et al., 2020). Nevertheless, these methodologies exhibit several drawbacks:
  1. (1)
@@
    The training of AI models is often data-intensive, with significant demands for an extended and tedious data collection through simulations or hardware experiments.
  2. (2)
@@
In recent years, large language models, such as GPT-4 (OpenAI, 2023), Palm (Anil et al., 2023) and LLaMa (Touvron et al., 2023) demonstrates superior performance in natural language understanding and generation. Furthermore, recent works introduce LLM-based autonomous agents which extend LLM’s capacity from simply performing reasoning and generating content to actions and control (Wang et al., 2024).
In our work, we investigate the potential of using LLM-based autonomous agents for downstream task automation in Power Electronics Systems. Specifically, we introduce LP-COMDA, a LLM-based Physics-Informed Power Converter Modulation Design Automation System, for users to produce high quality modulation design with minimal human supervision.
LP-COMDA first uses a LLM-based planner to process user requirements and generate a set of design specifications. Subsequently, the agent invokes a set of system design tools that consist of physics-informed surrogate models and optimization algorithms. In the design process, the tools iteratively derives optimal modulation parameters tailored to users’ design specifications. Finally, the optimal design parameters and performance metrics will be returned to and visualised for the users to achieve an explainable design process.
@@
  • We propose LP-COMDA, a LLM-based physics-informed autonomous agent that streamlines the power converter modulation design process. Comprehensive experiments shows strong performance of LP-COMDA, statistically outperforming the best baseline model with 63.2% lower error for low-data scenario, and 23.7% lower error for high-data scenario.
@@
  • Our user study with 20 experts demonstrates that the design time with LP-COMDA is over 33 times faster than conventional methods, showing superior efficiency in empirical practice with minimal human supervision.
Figure 1. An example of power converter application: the DAB (Dual Active Bridge) converter serves as a DC transformer of the DC-DC power grid, offering galvanic isolation and regulating power and voltage between DC buses. Modulating its switches directly affects the system operating performance, including power transfer efficiency, voltage regulation, and stability of the interconnected buses.
Figure 2. System architecture of LP-COMDA: an engineer provides design requirements to LP-COMDA via a chat interface connecting to its planner. Once the full requirements are determined, the planner coordinates and invokes tools from the tool set to iteratively generate the modulation design without human supervision. After the design is done, the planner displays the final results and explainable process on the chat interface.
2. Problem Statements
2.1. Power Converter Phase Shift Modulation
Within the renewable energy systems sector, power converters need to work in harmony with other components to ensure optimal functionality. Figure 1 depicts a classic application of a particular power converter, the dual active bridge (DAB) converter. This converter functions as a DC transformer, connecting various DC buses to facilitate the utilization or storage of power produced by RES (Lin et al., 2021). The phase shift modulation technique is employed to regulate the current flow and guarantee efficient power transmission. This is done by adjusting the timing of different power switch sets, represented as $S_{1}$ to $S_{8}$ in Figure 1. Table 1 details the distinct phase shift modulation strategies associated with the DAB converter, highlighting their respective degrees of freedom (DoF) and modifiable parameters (Hou and Li, 2019). It is worth noting that the adjustable parameter for $<S_{1},S_{5}>$ indicates the phase difference between $S_{1}$ and $S_{5}$.
@@ 3 $D_{1}$ $D_{o}$ $D_{2}$
Table 1. Phase shift modulation strategies with the number of degrees of freedom (DoF) from 1 to 3. $<S_{x},S_{y}>$ defines the phase shift between $S_{x}$ and $S_{y}$.
For designing the parameters of the modulation strategy, users are required to define three essential pieces of specification consistent with the application prerequisites:
  1. (1)
@@ 2.2. Problem Statements
3. Related work
3.1. LLM-based Autonomous Agents
The evolution of Large Language Models (LLMs) has unlocked immense potential for automating various processes in engineering and natural sciences through simple, plain-language queries. Numerous studies have explored LLM capability in reasoning and formal system applications, such as programming co-pilots (Ross et al., 2023; McNutt et al., 2023), multi-lingual tasks (Xu et al., 2024; Liu et al., 2024b), and mathematical problem-solving (Shakarian et al., 2023; Shi et al., 2024).
To address more complex industrial tasks, recent work proposed the use of LLM-based autonomous agents (LLM-agents), which extends the capability of LLMs (Wang et al., 2024). LLMs are deployed as the planning backbone to understand user request and invoke suitable tools to perform actions (Yao et al., 2024). Besides the planning module, LLM-agents also include other components, such as memory and tools. With advanced techniques such as Retrieval Augmented Generation (RAG) (Liu et al., 2024b; Xu et al., 2024), LLM-agents become powerful tools for solving complex tasks with little to no human instruction.
Recent works reported outstanding performance of LLM-agents in industrial applications. For instance, Cao and Lee harnessed LLMs and Phase-Step prompts for cross-domain robot task generation (Cao and Lee, 2023). Li et al. introduced an agentic framework to for strategic and interactive decision-making (Li et al., 2024b). LLM-based agents are found effective in other domain-specific tasks, such as spatial-temporal strategic planning (Li et al., 2024a; Liu et al., 2024a; Xiang et al., 2024), financial trading (Zhang et al., 2024b; Koa et al., 2024, 2024; Zhang et al., 2024a; Gao et al., 2023), and daily routine with computers and smart phones (Lai et al., 2024; Xing et al., 2024).
@@ 3.1. LLM-based Autonomous Agents
Figure 3. The proposed surrogate model consists of two physics-informed neural networks, namely, ModNet for switch-level modeling to learn the switching behaviors, and CirNet for system-level modeling to learn the circuit physics. The hierarchical structure enhances the overall accuracy of the power converter’s modeling of the complex behaviors of the switches.
3.2. Physics-Informed Neural Networks
@@
Conventional neural networks are commonly data-hungry and not explainable. By integrate physics principles into neural networks (Raissi, 2018), studies on Physics-Informed Neural Networks (PINNs) show promising results in overcoming such challenges.
@@
PINN encompasses various approaches to incorporate physics-informed (PI) components into the deep learning pipeline, such as loss functions, parameter initializations and the neural network architectures (Huang and Wang, 2022). As a result, the neural networks are able to learn latent features that are governed by physical laws.
@@
Recent studies also enhance PINN’s convergence, stability, and accuracy, yielding improvements (Kang et al., 2023; Yang et al., 2023; Wandel et al., 2022). PINN has found success in a wide range of applications. Some examples include predicting traffic conditions and modeling traffic flow (Shi et al., 2021; Ji et al., 2022), determining circuit parameters for power converter health monitoring (Zhao et al., 2022), computing power flow in electrical systems (Hu et al., 2020), modeling temperature dynamics in lakes (Jia et al., 2021), managing real-time reservoir gas production (Mudunuru et al., 2020), designing quantum circuits (Liu et al., 2021; Guo et al., 2019), and simulating turbulent flows in urban environments (Xiao et al., 2019). These applications highlight PINNs’ promising performance across many domains by infusing physics knowledge into deep learning.
4. Methodology
4.1. Agentic AI System
@@
Figure 2 presents the architecture of the proposed LP-COMDA agentic AI system. The front-end consists of a chat interface and API interaction with GPT-4 as reasoning engine, whereas the back-end consists of the agent’s tooling, i.e., an optimization algorithm and the PINN-based surrogate model. In terms of workflow, LP-COMDA first collects design specifications from the user and pass to an LLM-based reasoning engine (GPT-4) for planning. Subsequently, the agent invokes the back-end tooling, where the surrogate model produces the performance metrics and passes to the optimization algorithm searches for the optimal modulation parameters. The back-end will respond with a set of optimal design parameters to the front-end to display and visualise as charts for explainability and readability.
4.2. Physics-Informed Power Electronics Modeling
@@
(1) $\frac{\partial u(t)}{\partial t}=g(x(t);\theta)u(t)+h(x(t);\theta)$
@@
(2) $(\hat{v}_{p}(t_{k}),\hat{v}_{s}(t_{k}))=\operatorname{ModNet}(\hat{v}_{p}(t_{k-1}),\hat{v}_{s}(t_{k-1});\theta_{mod})$
@@
(3) $L\frac{\partial i_{L}(t)}{\partial t}=-R_{L}i_{L}(t)+v_{p}(t)-nv_{s}(t)$
@@
(4) $C_{1}\frac{\partial^{2}v_{C1}(t)}{\partial^{2}t}=$ $-\left(\frac{\partial s_{pri}(t)}{\partial t}-\frac{s_{pri}(t)R_{L}}{L}\right)i_{L}(t)+$
@@
(5) $C_{2}\frac{\partial^{2}v_{C2}(t)}{\partial^{2}t}=$ $n\left(\frac{\partial s_{sec}(t)}{\partial t}-\frac{s_{sec}(t)R_{L}}{L}\right)i_{L}(t)-$
@@
(6) $\hat{i}_{L}(t_{k+1})=\operatorname{CirNet}($ $\hat{i}_{L}(t_{k}),\operatorname{ModNet}($
@@
Circuit Performance Evaluation Based on CirNet: With the output of the two hierarchical PINNs, ModNet and CirNet, the power converter performance can be evaluated. With the key waveform inference results from the previous CirNet, a variety of electrical performance metrics can be measured, aiming to provide critical and comprehensive assessments for power electronics systems. As shown in Figure 3 Part (3), the difference of the maximum and minimum values of $i_{L}(t)$ gives the peak-to-peak current stress. The numerical integration of $[s_{1}(t)i_{L}(t)]^{2}$ provides the conduction loss evaluation for semiconductor switches. Soft switching performance, which is especially important for high-frequency scenarios like the electric vehicle charging, is analyzed by imposing constraints on $i_{L}(t)$ at the device commutation moment, where $s_{pri}(t)$ and $s_{sec}(t)$ are needed for synchronization. Other electrical performances can be gauged with the domain expertise in power electronics.
@@
5. Experiments
Figure 4. LP-COMDA with different structures of CirNet. The minimal MAE of 0.235 on validation set is reached when there are 2 hidden layers with 32 hidden neurons.
@@ 5. Experiments
LP-COMDA 19,857 2 layers, 32 hidden size
Table 2. Hyperparameter selection of LP-COMDA and benchmarks. The chosen hyperparameters yield the lowest loss on validation set.
5.1. Experiments Setup
We primarily aim to ascertain the accuracy of the ModNet and CirNet surrogate models, where heightened model accuracy in the core engine translates to superior system design performance. Therefore, we carried out several experiments to validate LP-COMDA’s model performance in a low-resource setting, i.e., using a tiny dataset with only 200 samples.
Data Preparation: In this work, we focus on a modeling task of time-series forecasting for the inductor current $i_{L}(t)$ of DAB converters under triple phase-shift (TPS) modulation as a representative example. Using a hardware experimental prototype, we acquired waveform data spanning 200 sequences. The configuration of the hardware prototype and the methodology employed for data acquisition are elucidated in reference (Lin et al., 2023b).
@@ 5.1. Experiments Setup
Baseline: To evaluate LP-COMDA’s design outcomes we use the following baselines: (1) the Bayesian Network (BN) (Pearl, 1985), (2) Support Vector Regression (SVR) (Smola and Schölkopf, 2004), (3) XGBoost (Chen et al., 2015), (4) Random Forest (RF) (Breiman, 2001), (5) Long-Short-Term-Memory (LSTM) (Hochreiter and Schmidhuber, 1997), (6) GRU net (Chung et al., 2014), (7) LN-GRU net (Ba et al., 2016), (8) Temporal Convolutional Net (TCN) (Lea et al., 2017), (9) GRU-Based Variational Autoencoder (GRU-VAE) (An and Cho, 2015), (10) Time-Series Transformer (TST) (Zerveas et al., 2021), (11) Time-Series Net Adapted from Vision Transformer (TSiT-Plus) (Oguiza, 2022), and (12) MiniRocket (Tan et al., 2022).
@@
Figure 5. Modeling performance of the top 5 benchmarks and the proposed LP-COMDA. The modelling results of LP-COMDA is closest to the measurement, demonstrating its outstanding modeling accuracy.
@@
Hyperparameter Search: For fair comparison, we conduct hyperparameter search on the validation set for LP-COMDA and all baseline algorithms. We perform grid search on important hyperparameters for each model. Figure 4 illustrates the MAEs of LP-COMDA with different hyperparameters, shedding light on the selected best structure. The final selected sets of hyperparameters are summarized in Table 2.
5.2. Experimental Results
Model Training Validation Test p-value
BN 2.152$\pm$0.0 2.436$\pm$0.0 2.515$\pm$0.0 6.81E-11*
SVR 1.984$\pm$0.0 2.313$\pm$0.0 2.286$\pm$0.0 1.16E-10*
XGBoost 2.299$\pm$0.0 2.108$\pm$0.0 2.792$\pm$0.0 3.83E-11*
RF 4.983$\pm$0.0 4.181$\pm$0.0 4.434$\pm$0.0 3.18E-12*
LSTM 1.164$\pm$0.151 1.380$\pm$0.042 1.560$\pm$0.095 7.23E-8*
GRU 1.287$\pm$0.174 1.336$\pm$0.101 1.604$\pm$0.064 6.35E-10*
LN-GRU 0.834$\pm$0.272 1.289$\pm$0.165 1.257$\pm$0.117 1.71E-6*
TCN 1.571$\pm$0.210 1.996$\pm$0.257 1.538$\pm$0.501 1.436E-3*
TST 0.668$\pm$0.062 0.750$\pm$0.028 0.666$\pm$0.063 1.47E-6*
TSiTPlus 0.707$\pm$0.158 1.106$\pm$0.12 1.302$\pm$0.078 3.46E-8*
MiniRocket 0.113$\pm$0.122 0.742$\pm$0.365 0.763$\pm$0.338 0.0131*
GRU-VAE 1.928$\pm$0.201 2.190$\pm$0.180 2.219$\pm$0.172 6.47E-7*
CirNet Only 0.121$\pm$0.014 0.282$\pm$0.014 0.310$\pm$0.013 1.442E-3*
LP-COMDA 0.140$\pm$0.054 0.235$\pm$0.025 0.245$\pm$0.029 N/A

### v2

# Physics-Informed Autonomous LLM Agents for Explainable Power Electronics Modulation Design

@@ Abstract
LLM-based autonomous agents have recently shown strong capabilities in solving complex industrial design tasks. However, in domains aiming for carbon neutrality and high-performance renewable energy systems, current AI-assisted design automation methods face critical challenges in explainability, scalability, and practical usability. To address these limitations, we introduce PHIA (Physics-Informed Autonomous Agent), an LLM-driven system that automates modulation design for power converters in Power Electronics Systems with minimal human intervention. In contrast to traditional pipeline-based methods, PHIA incorporates an LLM-based planning module that interactively acquires and verifies design requirements via a user-friendly chat interface. This planner collaborates with physics-informed simulation and optimization components to autonomously generate and iteratively refine modulation designs. The interactive interface also supports interpretability by providing textual explanations and visual outputs throughout the design process. Experimental results show that PHIA reduces standard mean absolute error by 63.2% compared to the second-best benchmark and accelerates the overall design process by over 33 times. A user study involving 20 domain experts further confirms PHIA’s superior design efficiency and usability, highlighting its potential to transform industrial design workflows in power electronics.
Introduction
@@
As global temperatures rise due to climate change, transitioning power generation systems toward carbon neutrality is increasingly urgent. A key component of this transition is the integration of renewable energy sources (RES) into the power grid. Power Electronics Systems (PES) are critical to this effort for three primary reasons (Lin 2022). Firstly, power converters within PES transform the direct current (DC) output from solar panels or wind turbines to alternating current (AC), facilitating the seamless integration of RES into the AC grid. Secondly, PES manages power during peak and off-peak periods, guiding the charge and discharge processes of energy storage units. Lastly, it upholds grid stability through functions like voltage regulation and reactive power compensation. Achieving the desired outcomes necessitates meticulous modulation of power converters in PES, which involves controlling their switching to manage output voltage or current.
@@
The escalating adoption of RES and the burgeoning scale of multi-resource power systems complicate the modulation design for power converters within PES. The diversity of energy sources linked to the power grid intensifies system complexity. As more converter subsystems interlink, the modulation design problem’s dimensionality surges. Consequently, identifying an optimized design solution through manual computation becomes infeasible given this complexity. As cities worldwide embrace multi-resource power systems, modulation designs must be customized, catering to particular geographies and application objectives. Oftentimes, these objectives can conflict, further complicating matters.
@@
The recent advancements in artificial intelligence (AI) and Large Language Models (LLMs) offer a promising solution to these modulation design challenges through automation. Emerging research has introduced AI-centric modulation design techniques. For instance, recent works explore the use of an LLM with in-context learning for modulation design  (Lin et al. 2024). Extreme gradient boosting algorithm (XGBoost) was utilised to construct the surrogate model for the triple phase shift modulation strategy for the dual active bridge (DAB) converter using training data sourced from simulation tools or hardware prototypes (Lin et al. 2023a). Subsequently, a differential evolution algorithm collaborates iteratively with XGBoost until optimal modulation parameters are identified. Likewise, the Q-learning algorithm, a conventional reinforcement learning technique, can be trained offline to derive optimized modulation parameters (Tang et al. 2020). While innovative, these methods suffer from several drawbacks:
  1. 1.
@@
    The training of these AI models is often data-intensive, with significant demands for an extended and tedious data collection through extensive simulations or hardware experiments.
  2. 2.
@@
In recent years, large language models, such as GPT-4 (OpenAI 2023), Palm (Anil et al. 2023) and LLaMa (Touvron et al. 2023) demonstrates superior performance in natural language understanding and generation. Furthermore, recent works introduce LLM-based autonomous agents which extend LLM’s capacity from simply performing reasoning and generating content to actions and control (Wang et al. 2024).
In our work, we investigate the potential of using LLM-based autonomous agents for downstream task automation in Power Electronics Systems. Specifically, we introduce PHIA, a Physics-Informed Autonomous Agent for power converter modulation design, enabling users to produce high quality modulation designs with minimal human supervision.
PHIA first uses a LLM-based planner to process user requirements and generate a set of design specifications. Subsequently, the agent invokes a set of system design tools that consist of physics-informed surrogate models and optimization algorithms. In the design process, the tools iteratively derives optimal modulation parameters tailored to users’ design specifications. Finally, the optimal design parameters and performance metrics will be returned to and visualised for the users to achieve an explainable design process.
@@
  • We propose PHIA, a physics-informed autonomous agent that streamlines the power converter modulation design process. Comprehensive experiments shows strong performance of PHIA, statistically outperforming the best baseline model with 63.2% lower error for low-data scenario, and 23.7% lower error for high-data scenario.
@@
  • Our user study with 20 experts demonstrates that the design time with PHIA is over 33 times faster than conventional methods, showing superior efficiency in empirical practice with minimal human supervision.
Figure 1: An example of power converter application: the DAB (Dual Active Bridge) converter serves as a DC transformer of the DC-DC power grid, offering galvanic isolation and regulating power and voltage between DC buses. Modulating its switches directly affects the system operating performance, including power transfer efficiency, voltage regulation, and stability of the interconnected buses.
Figure 2: System architecture of PHIA: an engineer provides design requirements to PHIA via a chat interface connecting to its planner. Once the full requirements are determined, the planner coordinates and invokes tools from the tool set to iteratively generate the modulation design without human supervision. After the design is done, the planner displays the final results and explainable process on the chat interface.
Problem Statements
Power Converter Phase Shift Modulation
Within the renewable energy systems sector, power converters need to work in coordination with other components to ensure optimal functionality. Figure 1 depicts a classic application of a particular power converter, the dual active bridge (DAB) converter. This converter functions as a DC transformer, connecting various DC buses to facilitate the utilization or storage of power produced by RES (Lin, Zhang, and Li 2021). The phase shift modulation technique is employed to regulate the current flow and guarantee efficient power transmission. This is done by adjusting the timing of different power switch sets, represented as $S_{1}$ to $S_{8}$ in Figure 1. Table 1 details the distinct phase shift modulation strategies associated with the DAB converter, highlighting their respective degrees of freedom (DoF) and modifiable parameters (Hou and Li 2019). It is worth noting that the adjustable parameter for $S_{1},S_{5}>$ indicates the phase difference between $S_{1}$ and $S_{5}$.
@@ 3 $D_{1}$ $D_{o}$ $D_{2}$
Table 1: Phase shift modulation strategies with the number of degrees of freedom (DoF) from 1 to 3. $S_{x},S_{y}>$ defines the phase shift between $S_{x}$ and $S_{y}$.
For designing the parameters of the modulation strategy, users are required to define three essential pieces of specification consistent with the application prerequisites, which are:
  1. 1.
@@
Figure 3: The proposed surrogate model consists of two physics-informed neural networks, namely, ModNet for switch-level modeling to learn the switching behaviors, and CirNet for system-level modeling to learn the circuit physics. The hierarchical structure enhances the overall accuracy of the power converter’s modeling of the complex behaviors of the switches.
Related work
LLM-based Autonomous Agents
The evolution of Large Language Models (LLMs) has unlocked immense potential for automating various processes in engineering and natural sciences through simple, plain-language queries. Numerous studies have explored LLM capability in reasoning and formal system applications, such as programming co-pilots (Ross et al. 2023; McNutt et al. 2023), multi-lingual tasks (Xu et al. 2024; Liu et al. 2024b), and mathematical problem-solving (Shakarian et al. 2023; Shi et al. 2024).
To address more complex industrial tasks, recent work proposed the use of LLM-based autonomous agents (LLM-agents), which extends the capability of LLMs (Wang et al. 2024). LLMs are deployed as the planning backbone to understand user request and invoke suitable tools to perform actions (Yao et al. 2024). Besides the planning module, LLM-agents also include other components, such as memory and tools. With advanced techniques such as Retrieval Augmented Generation (RAG) (Liu et al. 2024b; Xu et al. 2024), LLM-agents become powerful tools for solving complex tasks with little to no human instruction.
Recent works reported outstanding performance of LLM-agents in industrial applications. For instance, Cao and Lee harnessed LLMs and Phase-Step prompts for cross-domain robot task generation (Cao and Lee 2023). Li et al. introduced an agentic framework to for strategic and interactive decision-making (Li et al. 2024a). LLM-based agents are found effective in other domain-specific tasks, such as spatial-temporal strategic planning (Li et al. 2024b; Liu et al. 2024a; Xiang et al. 2024), e-commerce customer supports (Liu et al. 2024b, 2026, 2025) and financial trading (Zhang et al. 2024b; Koa et al. 2024; Zhang et al. 2024a; Gao, Wang, and Yang 2023).
Physics-Informed Neural Networks
@@
Conventional neural networks are commonly data-hungry and not explainable. By integrating physics principles into neural networks (Raissi 2018), studies on Physics-Informed Neural Networks (PINNs) show promising results in overcoming such challenges.
@@
PINN encompasses various approaches to incorporate physics-informed (PI) components into the deep learning pipeline, such as loss functions, parameter initializations and the neural network architectures (Huang and Wang 2022). As a result, the neural networks are able to learn latent features that are governed by physical laws.
@@
Recent studies also enhance PINN’s convergence, stability, and accuracy, yielding improvements (Kang et al. 2023; Yang, Qiu, and Fu 2023; Wandel et al. 2022). PINN has found success in a wide range of applications. Some examples include predicting traffic conditions and modeling traffic flow (Shi, Mo, and Di 2021; Ji et al. 2022), determining circuit parameters for power converter health monitoring (Zhao et al. 2022), computing power flow in electrical systems (Hu et al. 2020), modeling temperature dynamics in lakes (Jia et al. 2021), managing real-time reservoir gas production (Mudunuru et al. 2020), designing quantum circuits (Liu et al. 2021; Guo et al. 2019), and simulating turbulent flows in urban environments (Xiao et al. 2019). These applications highlight PINNs’ promising performance across many domains by infusing physics knowledge into deep learning.
Methodology
Agentic AI System
@@
Figure 2 presents the architecture of the proposed PHIA agentic AI system. The front-end consists of a chat interface and API interaction with GPT-4 as reasoning engine, whereas the back-end consists of the agent’s tooling, i.e., an optimization algorithm and the PINN-based surrogate model. In terms of workflow, PHIA first collects design specifications from the user and passes to an LLM-based reasoning engine (GPT-4) for planning. Subsequently, the agent invokes the back-end tooling, where the surrogate model produces the performance metrics and passes to the optimization algorithm searches for the optimal modulation parameters. The back-end will respond with a set of optimal design parameters to the front-end to display and visualise as charts for explainability and readability.
Physics-Informed Power Electronics Modeling
@@
$\frac{\partial u(t)}{\partial t}=g(x(t);\theta)u(t)+h(x(t);\theta)$ (1)
@@
$(\hat{v}_{p}(t_{k}),\hat{v}_{s}(t_{k}))=\operatorname{ModNet}(\hat{v}_{p}(t_{k-1}),\hat{v}_{s}(t_{k-1});\theta_{mod})$ (2)
@@
$L\frac{\partial i_{L}(t)}{\partial t}=-R_{L}i_{L}(t)+v_{p}(t)-nv_{s}(t)$ (3)
@@
$C_{1}\frac{\partial^{2}v_{C1}(t)}{\partial^{2}t}=$ $-\left(\frac{\partial s_{pri}(t)}{\partial t}-\frac{s_{pri}(t)R_{L}}{L}\right)i_{L}(t)+$ (4)
@@
$C_{2}\frac{\partial^{2}v_{C2}(t)}{\partial^{2}t}=$ $n\left(\frac{\partial s_{sec}(t)}{\partial t}-\frac{s_{sec}(t)R_{L}}{L}\right)i_{L}(t)$- (5)
@@
$\hat{i}_{L}(t_{k+1})=\operatorname{CirNet}($ $\hat{i}_{L}(t_{k}),\operatorname{ModNet}($ (6)
@@
Circuit Performance Evaluation Based on CirNet: With the output of the two hierarchical PINNs, ModNet and CirNet, the power converter performance can be evaluated. With the key waveform inference results from the previous CirNet, a variety of electrical performance metrics can be measured, aiming to provide critical and comprehensive assessments for power electronics systems. As shown in Figure 3 Part (3), the difference of the maximum and minimum values of $i_{L}(t)$ gives the peak-to-peak current stress. The numerical integration of $s_{1}(t)i_{L}(t)]^{2}$ provides the conduction loss evaluation for semiconductor switches. Soft switching performance, which is especially important for high-frequency scenarios like the electric vehicle charging, is analyzed by imposing constraints on $i_{L}(t)$ at the device commutation moment, where $s_{pri}(t)$ and $s_{sec}(t)$ are needed for synchronization. Other electrical performances can be gauged with the domain expertise in power electronics.
@@
Experiments
Figure 4: PHIA with different structures of CirNet. The minimal MAE of 0.235 on validation set is reached when there are 2 hidden layers with 32 hidden neurons.
@@
PHIA 19,857 2 layers, 32 hidden size
Table 2: Hyperparameter selection of PHIA and benchmarks. The chosen hyperparameters yield the lowest loss on validation set.
Experiments Setup
We primarily aim to ascertain the accuracy of the ModNet and CirNet surrogate models, where heightened model accuracy in the core engine translates to superior system design performance. Therefore, we carried out several experiments to validate PHIA’s model performance in a low-resource setting, i.e., using a tiny dataset with only 200 samples.
Data Preparation: In this work, we focus on a modeling task of time-series forecasting for the inductor current $i_{L}(t)$ of DAB converters under triple phase-shift (TPS) modulation as a representative example. Using a hardware experimental prototype, we acquired waveform data spanning 200 sequences. The configuration of the hardware prototype and the methodology employed for data acquisition are elucidated in reference (Lin et al. 2023b).
@@
Baseline: We evaluate PHIA’s design outcomes using: (1) the Bayesian Network (BN) (Pearl 1985), (2) Support Vector Regression (SVR) (Smola and Schölkopf 2004), (3) XGBoost (Chen et al. 2015), (4) Random Forest (RF) (Breiman 2001), (5) Long-Short-Term-Memory (LSTM) (Hochreiter and Schmidhuber 1997), (6) GRU net (Chung et al. 2014), (7) LN-GRU net (Ba, Kiros, and Hinton 2016), (8) Temporal Convolutional Net (TCN) (Lea et al. 2017), (9) GRU-Based Variational Autoencoder (GRU-VAE) (An and Cho 2015), (10) Time-Series Transformer (TST) (Zerveas et al. 2021), (11) Time-Series Net Adapted from Vision Transformer (TSiT-Plus) (Oguiza 2022), and (12) MiniRocket (Tan et al. 2022).
@@
Figure 5: Modeling performance of the top 5 benchmarks and the proposed PHIA. The modelling results of PHIA is closest to the measurement, demonstrating its outstanding modeling accuracy.
@@
Hyperparameter Search: For fair comparison, we conduct hyperparameter search on the validation set for PHIA and all baseline algorithms. We perform grid search on important hyperparameters for each model. Figure 4 illustrates the MAEs of PHIA with different hyperparameters, shedding light on the selected best structure. The final selected sets of hyperparameters are summarized in Table 2.
Experimental Results
Model
Training Validation Test p-value
[…]